OpenStudy (anonymous):

The cost of two coffees and four icecreams is 22 cents. The cost of three coffees and five icecreams is 30 cents. Value of a coffe is? Value of an icecream is?

OpenStudy (hexagon001):

let x be coffess and y be ice cream so 2x+4 =22 3x+5y=30 now solve these simulataneously

OpenStudy (anonymous):

Yes, but it isn't quite that easy. You try.

OpenStudy (anonymous):

c = coffee i = ice cream 2c + 4i = .22 ---> (3)2c + 4i = .22 3c + 5i = .30 --->(-2)3c + 5i = .30 -------------- 6c + 12i = .66 (result of multiplying by 3) -6c - 10i = - .60 (result of multiplying by -2) --------------add 0 + 2i = .06 ......i = .06/2 ......i = .03 now sub .03 in for i in either of the original equations 2c + 4i = .22 2c + 4(.03) = .22 2c + .12 = .22 2c = .22 - .12 2c = .10 c = .10/2 c = .05 check by subbing in known values into either of the original equations 3c + 5i = .30 2c + 4i = .22 3(.05) + 5(.03) = .30 2(.05) + 4(.03) = .22 .15 + .15 = .30 .10 + .12 = .22 .30 = .30 (correct) .22 = .22 (correct) Coffee costs 5 cents and ice cream costs 3 cents Do you understand what I just did ? Do you have any questions ?

OpenStudy (anonymous):

I do have a question. Why do we multiply by a negative 2?

OpenStudy (anonymous):

This is an elimination problem. That means you have to eliminate a variable in order to find the value of the other variable. I chose to eliminate the c variable. In order to do that, I had to make the c's opposite of each other. One has to be a negative 6 and one has to be a positive 6. If one is negative and one is positive, of course having the same coefficient, then they will cancel each other out. -6 and + 6 cancel(eliminate) each other. Thus, letting us solve for i. Does that make sense ?

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

your welcome :) Glad to help